# Quantum Turbulence Across Dimensions: Crossover from two- to three-dimension

### v1

@@ Abstract
We investigate the dynamic transition of quantum turbulence (QT) in a confined potential field as the system evolves from purely two-dimensional (2D) to quasi-two-dimensional, and ultimately to three-dimensional (3D), by fixing the lateral dimensions of the trapping box while varying its height. In the 2DQT, distinct Onsager vortex cluster formation and inverse energy cascade are observed, while 3DQT exhibits a direct energy cascade consistent with the Vinen turbulence decay rate, which display striking differences. By systematically altering the system height, we explore how dimensionality drives the differentiation of turbulence types and find that this transition is closely related to the excitation of Kelvin waves. Kelvin waves not only introduce additional dissipation mechanisms but also serve as mediators for direct energy transfer across scales. When the wavelength of the permitted Kelvin waves exceeds the critical size of vortex clusters, turbulence begins transitioning to 3D type, culminating in fully developed 3DQT at the characteristic scale. In the transitional region, we observe continuous variations in the decay rate and vortex cluster correlation functions.
@@
Turbulence is a ubiquitous nonlinear and complex phenomenon in nature, characterized by energy transfer across multiple scales and fundamentally driven by multiscale interactions [1, 2, 3]. The origins of turbulence can be traced back to the 19th century, when its chaotic nature was first observed experimentally, while theoretical descriptions primarily rely on the Navier-Stokes equations to capture its dynamic behavior. A key distinction between 2D and 3D Turbulence lies in the directionality of energy cascades: 2D turbulence exhibits an anomalous inverse energy cascade, where energy is transferred from small-scale vortices to larger scales, ultimately forming large-scale vortex clusters. This phenomenon, first revealed by Kraichnan and others through theoretical and numerical studies, arises due to the conservation of vorticity and enstrophy in 2D flows [4, 5]. In 2D turbulence, the enstrophy (the integral of the square of vorticity) cascades forward to smaller scales, while the conservation of total vorticity drives energy to accumulate at larger scales, leading to the formation of coherent vortex structures through the inverse cascade mechanism [6, 7]. In contrast, 3D turbulence exhibits a direct energy cascade, where energy is transferred from large scales to smaller scales and dissipated [8, 9]. Despite the significant differences between 2D and 3D turbulence, the mechanisms governing the transition between them and their dynamic characteristics under gradually relaxed geometric constraints remain incompletely understood.
@@
Quantum turbulence (QT), as a quantum counterpart of classical turbulence, provides a unique platform to study turbulence dynamics with unprecedented clarity and precision. Unlike classical turbulence, QT is composed of quantized vortices, which are topological defects characterized by a fixed circulation and a well-defined core structure [10, 11, 12]. This quantization not only makes quantized vortices highly identifiable and stable but also allows for direct experimental visualization and precise theoretical modeling. The discrete nature of these vortices eliminates the ambiguity associated with vorticity in classical systems, offering a cleaner and more tractable framework for studying turbulence.
In 3DQT, the dynamics are enriched by unique phenomena such as vortex reconnections and the excitation of Kelvin waves along vortex filaments. Vortex reconnections play a pivotal role in redistributing energy and generating cascades, while Kelvin waves mediate energy dissipation at smaller scales [13, 14, 15]. These mechanisms make 3DQT an ideal system to understanding energy dissipation and cascade processes that would otherwise be hidden within the chaotic, continuous vorticity fields of classical systems.
@@
In contrast, 2DQT exhibits distinct behaviors due to its alignment with the ideal point-vortex model. A key feature of 2DQT is the spontaneous formation of Onsager vortex clusters—large-scale, coherent structures that arise from the inverse energy cascade [16, 17, 18, 19, 20, 21, 22, 23, 19, 24, 25, 26, 27]. This clustering phenomenon is deeply rooted in statistical mechanics principles, where negative temperature states emerge as like-sign vortices aggregate into clusters while opposite-sign vortices annihilate [28]. The inverse energy cascade observed in 2DQT highlights the universality of this feature across classical and quantum systems, yet only in QT can the actual dynamics of individual vortices be directly observed and studied.
@@
In this Letter, we investigate the transition of QT from 2D to 3D, with the dimensional crossover driven by varying the system’s height. By imprinting the same number of quantized vortices in traps with different aspect ratios and allowing them to evolve freely, we observe how vortex dynamics exhibit characteristic features of different dimensional regimes. Through a comprehensive analysis of vortex clustering, decay dynamics, and energy spectra, we identify the key role played by Kelvin waves in mediating the transition between 2D and 3D QT.
@@
Here, $\psi(\mathbf{r},t)$ represents the condensate wavefunction, while $\hbar$, $M$, $\mu$, and $V_{\mathrm{trap}}$ are the reduced Planck constant, particle mass, chemical potential, and external trapping potential, respectively.
@@
Figure 1: Evolution of vortex dynamics for different potential trap heights $R_{z}$. The left column shows the evolution in the quasi-2D regime with $R_{z}=2$. The right column shows the evolution in the 3D regime with $R_{z}=10$. Movie can find in supplementary [29].
@@
Figure 2: (a) Vortex decay process and fitting line for $R_{z}=2$ (blue), $R_{z}=5$ (yellow) and $R_{z}=10$ (red) case. (b) Decay rate versus height.
As shown in Fig. 1, by using this boundary condition, we fixed the radius $R$ of the cylinder, thus simulating the transition from quasi-2D to 3D vortex dynamics by changing the height $R_{z}$. Initially, 100 positive and 100 negative vortices are randomly placed using the vortex imprinting method, and the system is allowed to undergo relaxation dynamics [27].
In the left half of Fig. 1, where a small height $R_{z}=2$ is selected, the vortex motion clearly exhibits two-dimensional characteristics. The vorticity remains along the $z$-direction, and vortex decay occurs only through vortex-antivortex pair annihilation. After a significant number of annihilations, two large vortex clusters, known as Onsager vortices, emerge, in agreement with experimental observations[19].
In contrast, the right half of Fig. 1 shows the system’s evolution at a larger height $R_{z}=10$. In this case, vortex lines quickly twist and entangle, forming the characteristic features of 3DQT. The vortices decay rapidly, and no coherent vortex clusters are formed, highlighting the suppression of Onsager vortex formation in the 3D regime.
@@
Due to the anisotropic nature of quasi-2DQT along the $z$-direction and the anisotropy of 3DQT, we consider the cluster correlation function along the $z$-direction and the dipole correlation function in the $x$ and $y$ directions. As shown in Fig. 3(a), during the evolution at $R_{z}=2$ in the quasi-2D regime, the cluster correlation function gradually increases and eventually stabilizes at a positive value, indicating the formation and stability of vortex clusters. When the height increases to $R_{z}=5$, although the result is less pronounced, stable vortex clusters still persist. However, at $R_{z}=10$, the cluster correlation function remains at zero, indicating the complete absence of vortex clusters.
@@
The cluster correlation function effectively characterizes this dynamic process, while the dipole correlation function provides further insight. As the height increases from small to large values, the dipole correlation function emerges and grows, corresponding to the formation of Kelvin waves. As shown in Fig. 4, the presence of Kelvin waves along vortex lines leads to the emergence of dipole-like structures observable from the side plane.
@@
In Fig. 3(b), we plot the height dependence of the cluster and dipole correlation functions, along with the corresponding phase diagram. As the height increases, the cluster correlation function begins to decrease at $R_{z}\geq 3$, which is consistent with the dimensional change of the decay rate, indicating that the inverse energy cascade of the two-dimensional features begins to be affected when the three-dimensional decay process occurs. At this point, the dynamics of Kelvin waves gradually start to dominate. When the height reaches $R_{z}\geq 6$, the dipole correlation function approaches its maximum, indicating the prevalence of Kelvin waves, and vortex clusters cease to form stably. Finally, for $R_{z}>7$, no vortex clusters are observed, and the system fully transitions to the 3DQT regime. This process highlights the critical role of Kelvin waves in driving the dimensional transition of QT, marking the shift from cluster-dominated dynamics in the 2D regime to reconnection- and wave-dominated dynamics in the 3D regime.
@@
Figure 5: Energy spectra for the pure 2D case (upper panel) and the 3D case (lower panel). The blue line represents the initial state, while the red line corresponds to the final stable state. In the pure 2D case, the three dotted vertical lines indicate the maximum energy peak scale, the minimum cluster scale, and the vortex scale, respectively.
@@
To further confirm the crucial role of Kelvin waves in the energy cascade, we calculate the quantum kinetic energy spectrum [36]:
[…]
@@
where $\mathcal{V}=\langle\psi\rangle\mathbf{u}$ and $\mathbf{u}$ is the superfluid velocity defined as $\mathbf{u}=\frac{i}{2}\left[\langle\psi^{\star}\rangle\nabla\langle\psi\rangle-\langle\psi\rangle\nabla\langle\psi\rangle^{\star}\right]/|\langle\psi\rangle|^{2}$.
@@
As shown in Fig. 6, in the purely 2DQT regime, the energy spectrum exhibits a clear inverse energy cascade at low wavenumbers, and the presence of a large-scale energy peak indicates the formation of vortex clusters [16]. Surprisingly, the scale of the energy peak is closely related to the critical height associated with the dimensional transition. Specifically, when the maximum possible Kelvin wavelength set by the system height satisfies
@@
the Kelvin waves start to introduce a directly energy cascade that interferes with the inverse cascade characteristic of 2DQT. This interference is reflected in the decrease of the decay rate and the cluster correlation function, as observed in Fig. 2 and Fig. 3. Eventually, when the system height exceeds $R_{z}>(2\pi/k)_{\mathrm{peak\ max}}\approx 7.9$, the inverse energy cascade completely vanishes, and the system fully transitions to the 3DQT regime. In this state, the energy spectrum no longer follows the Kolmogorov law with a $-5/3$ power law scaling [37]. This transition signifies that the dynamics of the Kelvin wave cascade dominate the system, effectively suppressing large-scale vortex clustering and driving the cascade processes characteristic of 3DQT.
In summary, by comparing the vortex relaxation dynamics at different potential trap heights, we have quantitatively studied, for the first time, the transition of QT behavior and energy cascades from two dimensions to three dimensions. The vortex decay rate, cluster correlation function, and energy spectrum demonstrate remarkable consistency throughout the transition. We conclude that the competition between the Kelvin wavelength and the characteristic vortex cluster size determines the dominance of 2D or 3D features. Only when the Kelvin wavelength exceeds the characteristic vortex cluster size do vortex reconnections and Kelvin-wave-mediated forward energy cascades begin to dominate. This leads to the disruption of large-scale structures typical of 2DQT and facilitates energy dissipation in the 3D regime, marking the onset of the dimensional transition. Kelvin waves, corresponding to the twisting of vortices in the transverse direction, play a crucial role in breaking vorticity conservation. This provides direct evidence that the conservation of vorticity is the key factor determining the direction of energy cascades. The breaking of this conservation allows for the forward cascade to emerge, ultimately driving the system into a fully developed 3DQT state. Experimentally, well-established techniques are already available to validate our results by systematically comparing the vortex dynamics at different potential trap heights[19].

### v2

@@ Abstract
We investigate the dynamic transition of quantum turbulence (QT) in a confined potential field, and by controlling the boundary aspect ratio, the system evolves from pure two-dimensional (2D) to quasi-two-dimensional, and finally to three-dimensional (3D). In the 2DQT, distinct Onsager vortex cluster formation and inverse energy cascade are observed, while 3DQT exhibits a direct energy cascade consistent with the Vinen turbulence decay rate, which display striking differences. By systematically altering the system height, we explore how dimensionality drives the differentiation of turbulence types and find that this transition is closely related to the excitation of Kelvin waves. Kelvin waves not only introduce additional dissipation mechanisms but also serve as mediators for direct energy transfer across scales. When the wavelength of the permitted Kelvin waves exceeds the critical size of vortex clusters, turbulence begins transitioning to 3D type, culminating in fully developed 3DQT at the characteristic scale. In the transitional region, we observe continuous variations in the decay rate and vortex cluster correlation functions.
@@
Turbulence is a ubiquitous nonlinear and complex phenomenon in nature, characterized by energy transfer across multiple scales and fundamentally driven by multiscale interactions [1, 2, 3]. The origins of turbulence can be traced back to the 19th century, when its chaotic nature was first observed experimentally, while theoretical descriptions primarily rely on the Navier-Stokes equations to capture its dynamic behavior. A key distinction between 2D and 3D Turbulence lies in the directionality of energy cascades: 2D turbulence exhibits an anomalous inverse energy cascade, where energy is transferred from small-scale vortices to larger scales, ultimately forming large-scale vortex clusters. This phenomenon, first revealed by Kraichnan and others through theoretical and numerical studies, arises due to the conservation of vorticity and enstrophy in 2D flows [4, 5]. In 2D turbulence, the enstrophy (the integral of the square of vorticity) cascades forward to smaller scales, while the conservation of total vorticity drives energy to accumulate at larger scales, leading to the formation of coherent vortex structures through the inverse cascade mechanism [6, 7]. In contrast, 3D turbulence exhibits a direct energy cascade, where energy is transferred from large scales to smaller scales and dissipated [8, 9]. Despite the significant differences between two- and three-dimensional turbulence, the mechanisms governing the transition between them and their dynamic characteristics under gradually relaxed geometric constraints remain elusive, though some studies have explored energy cascade changes [10, 11].
@@
Quantum turbulence (QT), which is widely observed in Bose-Einstein condensates[12, 13, 14] and superfluid 4He[15, 16], serves as the quantum counterpart of classical turbulence, offering a unique platform for studying turbulence dynamics. Unlike classical turbulence, QT consists of quantized vortices - topological defects characterized by fixed circulation and well-defined core structures [17, 18, 19]. This quantization not only makes vortices highly identifiable and stable but also enables direct experimental visualization and precise theoretical modeling, eliminating the ambiguity associated with vorticity in classical systems and providing a clearer, more tractable framework for turbulence research.
@@
In 3DQT, the dynamics are enriched by unique phenomena such as vortex reconnections and the excitation of Kelvin waves along vortex filaments. Vortex reconnections play a pivotal role in redistributing energy and generating cascades, while Kelvin waves mediate energy dissipation at smaller scales [20, 21, 22]. In contrast, 2DQT exhibits distinct behaviors due to its alignment with the ideal point-vortex model. A key feature of 2DQT is the spontaneous formation of Onsager vortex clusters—large-scale, coherent structures that arise from the inverse energy cascade [23, 24, 25, 26, 27, 28, 29, 30, 26, 31, 32, 33, 34]. This clustering phenomenon is deeply rooted in statistical mechanics principles, where negative temperature states emerge as like-sign vortices aggregate into clusters while opposite-sign vortices annihilate [35]. The inverse energy cascade observed in 2DQT highlights the universality of this feature across classical and quantum systems, yet only in QT can the actual dynamics of individual vortices be directly observed and studied. The consistency between quantum and classical turbulence in the dimension energy cascade provides a valuable framework for exploring the dimension crossing in the study of turbulence [36].
@@
In this Letter, we investigate the transition of QT from 2D to 3D, with the dimensional crossover driven by varying the system’s aspect ratio. By imprinting the same number of quantized vortices in traps with different aspect ratios and allowing them to evolve freely, we observe how vortex dynamics exhibit characteristic features of different dimensional regimes. Through a comprehensive analysis of vortex clustering, decay dynamics, and energy spectra, we identify the key role played by Kelvin waves in mediating the transition between 2D and 3D QT.
@@
Here, $\psi(\mathbf{r},t)$ represents the condensate wavefunction, while $\hbar$, $M$, $g$, $\mu$, and $V_{\mathrm{trap}}$ are the reduced Planck constant, particle mass, coupling constant, chemical potential, and external trapping potential, respectively.
Figure 1: Density evolution of vortex dynamics for different potential trap heights $R_{z}$. The left column shows the evolution in the quasi-2D regime with $R_{z}=2$, the final figure presents a cross-sectional view after the formation of stable vortex clusters, where positive and negative vortices are represented in green and purple, respectively. The right column shows the evolution in the 3D regime with $R_{z}=10$. Movie can find in supplementary [37].
@@
Figure 2: (a) Vortex decay process (solid) and fitting line (dash) for $R_{z}=2$ (blue), $R_{z}=5$ (yellow) and $R_{z}=10$ (red) case. (b) Decay rate versus height.
As shown in Fig. 1, by using this boundary condition, we fixed the radius $R=30$ of the cylinder, thus simulating the transition from quasi-2D to 3D vortex dynamics by changing the height $R_{z}$. Initially, 100 positive and 100 negative vortices are randomly placed using the vortex imprinting method, and the system is allowed to undergo relaxation dynamics [34].
In the left half of Fig. 1, where a small height $R_{z}=2$ is selected, the turbulence dynamic clearly exhibits two-dimensional characteristics. Vortices can be approximated as point vortices, and vortex decay occurs only through vortex-antivortex annihilation. After extensive annihilation, two large vortex clusters, known as Onsager vortices, emerge and keep stable, consistent with experimental observations [26]. The corresponding Movie can be found in the supplementary[37].
In contrast, the right half of Fig. 1 shows the system’s evolution at a larger height $R_{z}=10$. In this case, vortex lines quickly twist and entangle, forming the characteristic features of 3DQT. Driven by vortex reconnections and Kelvin waves, vortex decay occurs rapidly, preventing the formation of coherent vortex clusters, highlighting the suppression of Onsager vortex formation in the 3D regime.
@@
Due to the anisotropic nature of quasi-2D QT along the $z$-direction and the isotropy of 3DQT, we consider the cluster correlation function along the $z$-direction and the dipole correlation function in the $x$ and $y$ directions. As shown in Fig. 3(a), during the evolution at $R_{z}=2$ in the quasi-2D regime, the cluster correlation function gradually increases and eventually stabilizes at a large positive value $C_{4z}\approx 0.5$, indicating the formation and stability of Onsager vortex. When the height increases to $R_{z}=5$, although the result is less pronounced, stable vortex clusters still persist. However, at $R_{z}=10$, the cluster correlation function is always zero, indicating the complete absence of vortex clusters.
@@
The clustering correlation function effectively characterizes the dynamic evolution of QT, while the dipole correlation function provides further insights. As the height increases from small to large, the initial values of the dipole correlation function in the $x$ and $y$ directions emerge and grow, which corresponds to the presence of Kelvin waves. This can be explained using Fig. 4, where the existence of Kelvin waves leads to the appearance of dipole structures observable in the side cross-section of the vortex line.
@@
In Fig. 3(b), we plot the height dependence of the cluster and dipole correlation functions, along with the corresponding phase diagram. As the height increases, the cluster correlation function begins to decrease at $R_{z}\geq 3$, which is consistent with the dimensional change of the decay rate, indicating that the inverse energy cascade of the two-dimensional features begins to be affected at the same time as the three-dimensional decay process occurs. At this stage, Kelvin wave dynamics gradually become dominant. When the height reaches $R_{z}\geq 6$, tthe dipole correlation function approaches its maximum, signifying that the initial Kelvin wave density has peaked, after which vortex clusters lose their stability. Finally, for $R_{z}>7$, no vortex clusters are observed, the system fully transitions to the 3DQT regime. This process highlights the critical role of Kelvin waves in driving the dimensional transition of QT, marking the shift from cluster-dominated dynamics in the 2D regime to reconnection- and wave-dominated dynamics in the 3D regime.
@@
Figure 5: Energy spectra for the pure 2D case (upper panel) and the 3D case (lower panel). The blue line represents the initial state, while the red line corresponds to the final stable state. In the pure 2D case, the three dotted vertical lines represent the characteristic scales of the system: the maximum energy peak scale, the minimum cluster scale, and the vortex scale. The diagonal dashed lines correspond to the Kolmogorov spectrum.
@@
To further confirm the crucial role of Kelvin waves in the energy cascade, we calculate the quantum kinetic energy spectrum for the incompressible part [44, 45]:
$E_{k}^{i}(k)=\frac{1}{2}\int_{k\leq|\mathbf{k}^{\prime}|<k+\Delta k}\left|\mathbf{V}^{i}(\mathbf{k}^{\prime})\right|^{2}d^{3}k^{\prime}$ (6)
@@
where the incompressible field in Fourier space is given by $\mathbf{V}^{i}(\mathbf{k})=\mathcal{F}[\mathbf{V}]-(\mathbf{k}\cdot\mathcal{F}[\mathbf{V}])\mathbf{k}/|\mathbf{k}|^{2}$, here $\mathbf{V}=\psi\mathbf{u}$ and the superfluid velocity $\mathbf{u}$ defined as $\mathbf{u}=\frac{i}{2}(\psi^{*}\nabla\psi-\psi\nabla\psi^{*})/|\psi|^{2}$.
@@
As shown in Fig. 5, in the pure 2D QT state, the energy spectrum exhibits a clear inverse energy cascade at low wavenumbers. Additionally, an obvious energy peak emerges above the Kolmogorov law in the inertial range, which corresponds to the formation of vortex clusters, indicating the accumulation of energy at low wave numbers [23]. Surprisingly, the scale of the energy peak is closely related to the critical height associated with the dimensional transition. Specifically, when the maximum possible Kelvin wavelength allowed by the system height satisfies
@@
the Kelvin waves start to introduce a directly energy cascade that interferes with the inverse cascade characteristic of 2DQT. This interference is reflected in the decrease of the decay rate and the cluster correlation function, as observed in Fig. 2 and Fig. 3. Eventually, when the system height exceeds $R_{z}>(2\pi/k)_{\mathrm{peak\ max}}\approx 7.9$, the inverse energy cascade completely vanishes, and no vortex clusters form, indicating a full transition to the 3DQT regime. This transition signifies that the dynamics of the Kelvin wave cascade dominate the system, effectively suppressing large-scale vortex clustering and driving the cascade processes characteristic of 3DQT.
In summary, by comparing the vortex relaxation dynamics at different potential trap heights, we have quantitatively studied, for the first time, the transition of QT behavior and energy cascades from two dimensions to three dimensions. The vortex decay rate, cluster correlation function, and energy spectrum demonstrate remarkable consistency throughout the transition. We conclude that the competition between the Kelvin wavelength and the characteristic vortex cluster size determines the dominance of 2D or 3D features. Only when the Kelvin wavelength exceeds the characteristic vortex cluster size do vortex reconnections and Kelvin-wave-mediated forward energy cascades begin to dominate. This leads to the disruption of large-scale structures typical of 2DQT and facilitates energy dissipation in the 3D regime, marking the onset of the dimensional transition. Kelvin waves, corresponding to the twisting of vortices in the transverse direction, play a crucial role in breaking vorticity conservation. This provides direct evidence that the conservation of vorticity is the key factor determining the direction of energy cascades. The breaking of this conservation allows for the directly cascade to emerge, ultimately driving the system into a fully developed 3DQT state. Experimentally, well-established techniques are already available to validate our results by systematically comparing the vortex dynamics at different potential trap heights[26].